OpenStudy (anonymous):

How would I evaluate this: Log8 1/3

OpenStudy (anonymous):

\[\log_{b}(x) = {\log_{k}(x) \over \log_{k}(b)} \]

OpenStudy (anonymous):

log8(1/3)=log8(1)-log8(3) then log8(1)-log8(3)=log(1)/log(8) - log(3)/log(8)

OpenStudy (anonymous):

Thanks!